I have a problem figuring out this question can you help?
Find the ordered pair solution of y= 2/3 x -2 corresponding to x=3. Thank You
isbn 0618372628

Answer by Nate(3500)

y = (2/3)x - 2 since (x = 3), we can plug
y = 2 - 2 = 0
P(3,0)